# Is A Picture Worth A Thousand Words? Delving Into Spatial Reasoning for Vision Language Models

### v1

@@ Abstract
Large language models (LLMs) and vision-language models (VLMs) have demonstrated remarkable performance across a wide range of tasks and domains. Despite this promise, spatial understanding and reasoning—a fundamental component of human cognition—remains under-explored. We develop novel benchmarks that cover diverse aspects of spatial reasoning such as relationship understanding, navigation, and counting. We conduct a comprehensive evaluation of competitive language and vision-language models. Our findings reveal several counter-intuitive insights that have been overlooked in the literature: (1) Spatial reasoning poses significant challenges where competitive models can fall behind random guessing; (2) Despite additional visual input, VLMs often under-perform compared to their LLM counterparts; (3) When both textual and visual information is available, multi-modal language models become less reliant on visual information if sufficient textual clues are provided. Additionally, we demonstrate that leveraging redundancy between vision and text can significantly enhance model performance. We hope our study will inform the development of multimodal models to improve spatial intelligence and further close the gap with human intelligence.
@@ 1 Introduction
Visual understanding and reasoning—an intrinsic part of human perception and cognitive ability—has been largely under-explored when it comes to VLMs. In fact, it is often argued that the visual sense is the dominant sense in people, yet when it comes to current models, it seems quite secondary. Spatial reasoning, in particular, is fundamental to everyday human activities such as navigating environments, understanding maps, and manipulating objects. It encompasses skills that are crucial for both survival and higher-order cognition, including the ability to navigate through space, recognize patterns, and deduce relationships from spatial configurations.
In this paper, we develop three novel, visual-question answering (VQA) style benchmarks (Spatial-Map, Maze-Nav, and Spatial-Grid, details in Section 3.1) to explore the performance of VLMs on diverse aspects of spatial reasoning, including relationship, navigation, position understanding, and object counting. Humans excel at such tasks, making them essential capabilities for intelligent systems to emulate for safe and effective deployment in the real world.
@@
We conduct a systematic and comprehensive evaluation on a wide range of open-source and proprietary LLMs and VLMs. We perform in-depth analysis and unveil several novel and surprising results that challenge the current understanding of how these models handle spatial information:
@@
We surmise that the limitations in spatial understanding of VLMs arise from the overly simplistic treatment of images in the current VLM architectures and training pipelines. We believe that the contributions in this paper will drive changes in how these models are constructed so that they can unlock greater spatial reasoning capabilities, bridging the gap to human-like intelligence.
2 Related Work
Large language models. The generative large language models (LLMs) have been successful in recent years, which are based on the transformer architecture introduced by [61]. These LLMs, such as BERT [14], OPT [70], PaLM [12], Gemma [56], Mistral [25], GPT-3 [10], GPT-4 [2], Claude3 [5], Grok-1 [63], Llama [58], Llama2 [59], Llama3 [3] and so on, have profoundly changed the world due to their emergent ability, e.g., in-context learning [43, 45, 53], compositional ability [16, 19, 64] and adaptation ability [51, 52, 65]. LLMs have been involved in human activities, such as finance [31], bio-informatics [57], law [54], education [27], coding [23], and even creative writing [2] such as top AI conference reviews [33].
Large vision-language models and multi-modality. The success of LLMs has propelled the adoption of the transformer architecture within the computer vision community, such as ViT [15], Beit [9], CLIP [48], MAE [20], Swin [39, 40] and DiT [47]. Building on the capabilities of powerful LLMs, multi-modal language models (MLLMs) such as Flamingo [4], LLaMA-Adapter [18, 68], LLava [34, 36], stable-diffusion [49], BLIP [29, 30], MiniGPT-4 [71], Qwen [7, 8], Gemini [55], MM1 [42] have significantly expanded the scope of tasks that can be addressed. These models adeptly handle inputs from diverse modalities and have demonstrated remarkable performance on a wide range of tasks such as mathematical reasoning [41], visual question answering [6, 37], image-text retrieval [67], and visual reasoning [26].
@@ 3.1 Dataset Setup
To evaluate the spatial reasoning abilities of LLMs and VLMs, we construct three diverse tasks including spatial relationships, navigation, position understanding, and counting. To systematically study the impact of modality, we design three types of input prompts for each task: (1) Text-only: the input is purely textual and contains all necessary information for a person to answer the questions. (2) Vision-only: the input consists solely of an image, which provides sufficient details for a person to easily answer, a format also referred to as Visual Question Answering (VQA) in the literature. (3) Vision-text: the input includes both an image and its textual representation with detailed descriptions, rendering the information in both modalities redundant. We evaluate language models using Text-only inputs and multimodal language models using Text-only, Vision-only, and Vision-text inputs on the same set of questions. We opt for synthetic data in this work. Synthetic data offers several compelling advantages: (1) Avoidance of data leakage—since large language models are pre-trained on web-scale data, it is crucial to ensure that the test data have not been seen during training; (2) configurability—being fully configurable allows for controlled experiments and can be extended to additional tasks; (3) scalability—it is easy to scale the number of test samples, which enhances the statistical significance of our results.
Figure 1: Illustration of the Spatial-Map task, which simulates a map with multiple locations. To investigate the impact of modality, we consider three input formats: Text-only, Vision-only, and Vision-text. We evaluate language models (w. Text-only input) and multimodal language models (w. Vision-only and Vision-text inputs) on the same set of questions.
@@
Maze-Nav. Navigation through complex spaces is essential for intelligent systems. To evaluate such abilities, we have developed a maze-like dataset named Maze-Nav. Visually, each sample can be represented as colored blocks where different colors signify distinct elements: a green block marks the starting point (S), a red block indicates the exit (E), black blocks represent impassable walls, white blocks denote navigable paths, and blue blocks trace the path from S to E. The objective is to navigate from S to E following the blue path, with movement permitted in the four cardinal directions (up, down, left, right). Alternatively, each input can be depicted in textual format using ASCII code. An example is illustrated in Figure 2, featuring Text-only, Vision-only, and Vision-text inputs. We construct this task based on an open-sourced library [24]. The questions asked include counting the number of turns from S to E and determining the spatial relationship between S and E. While such questions are easy for humans, we will show in Section 4 that they present significant challenges for modern multimodal language models.
@@
Figure 2: Illustration of the Maze-Nav task, which evaluates the model’s ability to navigate from the starting point (S) to the exit (E). We consider three input formats: Text-only, Vision-only, and Vision-text. We evaluate language models (w. Text-only input) and multimodal language models (w. Vision-only and Vision-text inputs) on the same set of questions.
Figure 3: Illustration of the Spatial-Grid task, which evaluates the model’s spatial reasoning ability in a rigid grid structure. We consider three input formats: Text-only, Vision-only, and Vision-text. We evaluate language models (w. Text-only input) and multimodal language models (w. Vision-only and Vision-text inputs) on the same set of questions.
@@ 3.2 Models
We consider a wide range of competitive open-sourced language models with different scales, including Phi2-2.7B [32], the LLaMA family of models (LLaMA-2-7B, LLaMA-2-13B, and LLaMA-3-8B) [60], Mistral-7B [25], the Vicuna family (Vicuna-7B-1.5 and Vicuna-13B-1.5) [11], and Nous-Hermes-2-Yi-34B. For multimodal language models, we consider the Bunny family (Bunny-Phi-2-SigLIP, Bunny-Phi-1.5-SigLIP, Bunny-Phi-2-EVA, and Bunny-Phi-1.5-EVA) [21], CogVLM [62], CogAgent [22], InstructBLIP family (InstructBLIP-Vicuna-7B and InstructBLIP-Vicuna-13B) [13], and LLaVA family (LLaVA-1.6-Mistral-7B, LLaVA-1.6-Vicuna-7B, LLaVA-1.6-Vicuna-13B, and LLaVA-1.6-34B)  [35]. We also evaluate the proprietary models: Open AI’s GPT-4V, GPT-4O, GPT-4, Google Gemini Pro 1.0 and Anthropic Claude 3 Opus.
Evaluation. By default, 3 questions (termed Q1 to Q3) are associated with each sample in each task. As each question contains four options, we use accuracy as the main evaluation metric. The same user prompt is appended at the end of each question: First, provide a concise answer in one sentence. Then, elaborate on the reasoning behind your answer in a detailed, step-by-step explanation. For each model, we adopt the default configurations and decoding strategies, e.g., $\mathsf{argmax}$ for deterministic decoding and top-$p$ for non-deterministic decoding. For non-deterministic decoding, the results are averaged over three independent runs for open-sourced models. For proprietary models, due to their limited availability and increased compute time and cost, we only used one run.
@@ 4 Main Results and Analysis
Spatial reasoning remains surprisingly challenging. The evaluation results on open-sourced models on Spatial-Map, Maze-Nav, and Spatial-Grid are shown in Figure 4. For each task, the reported accuracy is averaged over all questions. For vision-language models, we choose the Vision-only input format, commonly used in Visual Question Answering (VQA). We use a dashed red line in each figure to indicate the expected accuracy if answering by random guessing. Our findings reveal several notable insights: (1) Vision-only inputs: Despite the ease with which humans handle these tasks, most competitive multimodal language models perform at levels comparable to or barely above random guessing. (2) Text-only inputs: The textual representation includes essential spatial information. However, such input format generally does not enhance the spatial reasoning capabilities of competitive language models. An exception occurs in the Spatial-Grid task, where Llama-3 achieves an accuracy of 71.9%, followed by Mistral-7B-Instruct at 62.1%, both notably surpassing random guessing. Despite these successes, the performance of these models still lags significantly behind human levels. These results underscore the need for further development of techniques tailored to spatial understanding and reasoning.
Figure 4: Performance overview on spatial reasoning tasks. We report the accuracy averaged over all questions. We consider the VQA (Vision-only) format for vision-language models. The dashed red line denotes the expected accuracy of random guessing. For Spatial-Map and Maze-Nav tasks, only a few models outperform random guessing by a notable margin.
Figure 5: TQA vs. VQA on spatial reasoning tasks. Each vertex on the spider plot represents the Avg Acc of a (VLM, LLM) pair with the same language backbone, i.e., LLM v.s. VLM further finetuned on that. VLMs are depicted in red, and LLMs in blue. We can see that VLMs rarely enhance the performance compared to their LLM counterparts.
The impact of input modality. To investigate the impact of modality, we compare the performance of a large language model (LLM) and a vision-language model (VLM) with the same language backbone. We consider the VQA (Vision-only) format for vision-language models. The results are shown in Figure 5. Each vertex on the spider plot represents the average accuracy of a (VLM, LLM) pair. We observe that on Spatial-Map and Spatial-Grid, the majority of VLMs yield worse performance compared to their LLM counterpart, despite having an additional visual encoder. For example, on Spatial-Grid, Mixtral-7B achieves an average accuracy of 62.1%, while LLaVA-v1.6-Mistral-7B only yields an accuracy of 47.1% (15% $\downarrow$). Detailed results can be seen in Appendix D.
@@ 5.1 Seeing Without Understanding: The Blindness of Multimodal Language Models
To better understand how multimodal language models process visual information, we conduct a series of controlled experiments in the VTQA (Vision-text input) setting. For each sample, we replace the original image input (that matches the textual description) with either: (1) No Image: only keep the textual input without the image input, (2) Noise Image: a Gaussian noise image irrelevant to the task, and (3) Random Image: a random image from the dataset that does not match the textual description, as shown in Figure 6.
@@
Noise image can improve the performance. We replace the Original Image with a Noise Image while retaining the original textual description. The results are shown in Figure 8. Consistent with the findings in Original Image vs. No Image, using a noise image also improves the performance across various VLM architectures. For example, the accuracy of LLaVA-1.6-Vicuna-13B increases by 6.5% on the Maze-Nav task when the noise image is used as opposed to the original image. In contrast to the ’No Image’ setting, noise images provide limited visual cues. Nonetheless, the model tends to prioritize the textual information, especially when visual cues are not pertinent to the task.
@@
Figure 7: Original Image vs. No Image in VTQA. VLMs exhibit improved performance in spatial reasoning tasks when visual input is absent.
@@ 5.2 Leveraging Redundancy in Multimodal Inputs
Multimodal language models offer considerable versatility in handling multimodal inputs. While the visual input alone often provides sufficient details for humans to address spatial reasoning tasks with ease, we propose that VLMs significantly benefit from the inclusion of textual descriptions alongside visual data, even if this introduces substantial redundancy. We verify this hypothesis by comparing VQA (Vision-only input) and VTQA (Vision-text input) across diverse VLM architectures. The results are shown in Figure 10, where each vertex on the spider plot represents the average accuracy of a (Vision-only, Vision-text) pair based on the same VLM. For Spatial-Map and Spatial-Grid, we can clearly see that having the additional textual input (VTQA) enhances the performance compared to only using images (VQA) across different VLM architectures. This suggests that textual inputs improve the accuracy of spatial reasoning in VLMs. We further compare TQA and VTQA in Appendix C. Detailed results are included in Appendix D.
@@
As many recent benchmarks have shown that proprietary models generally outperform open-sourced models, it is important to understand if our observed trends hold with proprietary models. The performance of several top proprietary models (GPT-4, GPT-4V, GPT-4O, Gemini Pro 1.0, and Claude 3 Opus) are shown in Figure 12. We have the following salient observations: (1) A significant performance gap exists between SoTA open-sourced models and proprietary models, as expected. Furthermore, with both Text-only and Vision-text formats, GPT-4V and GPT-4O significantly outperform random guessing across all tasks. For instance, in the Vision-text format, GPT-4O achieves an accuracy of 0.989 on Spatial-Grid (Table 2). (2) Yet, the trends we observed with open source models hold, VQA consistently under-performs compared to TQA and VTQA, for example, GPT-4V’s performance improves by 25.6% on Spatial-Grid when switching from Vision-only to Vision-text input; and again no clear winner between TQA and VTQA (see Appendix C for further details), showing that proprietary models, even the new GPT-4O model, still do not appear to fully leverage visual inputs.

### v2

@@ Abstract
Large language models (LLMs) and vision-language models (VLMs) have demonstrated remarkable performance across a wide range of tasks and domains. Despite this promise, spatial understanding and reasoning—a fundamental component of human cognition—remains under-explored. We propose SpatialEval, a novel benchmark that covers diverse aspects of spatial reasoning such as relationship understanding, navigation, and counting. We conduct a comprehensive evaluation of competitive language and vision-language models. Our findings reveal several counter-intuitive insights that have been overlooked in the literature: (1) Spatial reasoning poses significant challenges where competitive models can fall behind random guessing; (2) Despite additional visual input, VLMs often under-perform compared to their LLM counterparts; (3) When both textual and visual information is available, multi-modal language models become less reliant on visual information if sufficient textual clues are provided. Additionally, we demonstrate that leveraging redundancy between vision and text can significantly enhance model performance. We hope our study will inform the development of multimodal models to improve spatial intelligence and further close the gap with human intelligence. Our code is available at https://github.com/jiayuww/SpatialEval.
@@ 1 Introduction
Visual understanding and reasoning—an intrinsic part of human perception and cognitive ability—have been largely under-explored when it comes to LLMs and VLMs. In fact, it is often argued that the visual sense is the dominant sense in people, yet when it comes to current models, it seems quite secondary. Spatial reasoning, in particular, is fundamental to everyday human activities such as navigating environments, understanding maps, and manipulating objects. It encompasses skills that are crucial for both survival and higher-order cognition, including the ability to navigate through space, recognize patterns, and deduce relationships from spatial configurations.
In this paper, we propose SpatialEval, a novel benchmark containing four tasks (Spatial-Map, Maze-Nav, Spatial-Grid, and Spatial-Real, Section 2.1) to explore the performance of LLMs and VLMs on diverse aspects of spatial reasoning, including relationship, navigation, position understanding, and object counting. Humans excel at such tasks, making them essential capabilities for intelligent systems to emulate for safe and effective deployment in the real world.
@@
We conduct a systematic and comprehensive evaluation of a wide range of open-source and proprietary LLMs and VLMs. We perform in-depth analysis and unveil several novel and surprising results that challenge the current understanding of how these models process spatial information:
@@
We surmise the limitations in VLMs’ spatial understanding stem from the overly simplistic handling of visual information in current architectures and training pipelines. We believe the contributions of this paper will drive changes in model design, accelerating improvements that could unlock more robust spatial reasoning capabilities, and help bridge the gap toward human-like intelligence.
@@ 2.1 Dataset Setup
Figure 1: Illustration of the Spatial-Map task, which simulates a map with multiple locations. To investigate the impact of modality, we consider three input formats: Text-only, Vision-only, and Vision-text. We evaluate language models (w. TQA input) and vision-language models (w. VQA and VTQA inputs) on the same set of questions.
To evaluate the spatial reasoning abilities of LLMs and VLMs, we construct four diverse tasks including spatial relationships, navigation, position understanding, and counting. To systematically study the impact of modality, we design three types of input formats for each task: (1) TQA (Text-only): the input is purely textual and contains all necessary information for a person to answer the questions. (2) VQA (Vision-only): the input consists solely of an image, which provides sufficient details for a person to easily answer, a format also referred to as Visual Question Answering (VQA) in the literature. (3) VTQA (Vision-text): the input includes both an image and its textual representation with detailed descriptions, rendering the information in both modalities redundant. We evaluate LLMs using Text-only inputs and VLMs using Text-only, Vision-only, and Vision-text inputs on the same set of questions (Table 1). The synthetic tasks are curated based on the following key guidelines: (1) Avoidance of data leakage—since LLMs are pre-trained on web-scale data, it is crucial to ensure that the test data have not been seen during training; (2) configurability—being configurable allows for controlled experiments and extends easily to additional tasks; (3) scalability—the ability to scale the number of test samples enhances the statistical significance of experimental results.
@@
Figure 2: Illustration of the Maze-Nav task, which evaluates the model’s ability to navigate from the starting point (S) to the exit (E).
Maze-Nav. Navigation through complex spaces is essential for intelligent systems. To evaluate such abilities, we have developed a maze-like dataset named Maze-Nav. Visually, each sample can be represented as colored blocks where different colors signify distinct elements: a green block marks the starting point (S), a red block indicates the exit (E), black blocks represent impassable walls, white blocks denote navigable paths, and blue blocks trace the path from S to E. The objective is to navigate from S to E following the blue path, with movement permitted in the four cardinal directions (up, down, left, right). Alternatively, each input can be depicted in a textual format using ASCII code. An example is illustrated in Figure 2, featuring Text-only, Vision-only, and Vision-text inputs. We construct this task based on an open-sourced library [26]. The questions asked include counting the number of turns from S to E and determining the spatial relationship between S and E. While such questions are easy for humans, we will show in Section 3 that they still pose significant challenges for modern multimodal language models.
Figure 3: Illustration of the Spatial-Grid task, which evaluates the model’s spatial reasoning ability in a rigid grid structure.
@@
Figure 4: Illustration of the Spatial-Real task, which is built on real images with long captions, featuring detailed descriptions averaging over 1,000 words per image.
Spatial-Real. To extend the evaluation of spatial reasoning beyond synthetic environments, we introduce Spatial-Real, a task built on the Densely Captioned Images (DCI) dataset [67], where each image has a detailed caption with more than 1,000 words on average. As DCI does not contain questions, we curate multiple-choice questions regarding spatial reasoning (object counting, relation, and position understanding) and annotate the answers. An example is shown in Figure 4. We provide detailed analysis on Spatial-Real in Appendix E.
@@ 2.2 Models
We consider a wide range of competitive open-source language models with different scales, including Phi2-2.7B [35], the LLaMA family (LLaMA-2-7B, LLaMA-2-13B, and LLaMA-3-8B) [65], Mistral-7B [27], the Vicuna family (Vicuna-7B-1.5 and Vicuna-13B-1.5) [11], and Nous-Hermes-2-Yi-34B. For multimodal language models, we consider the Bunny family (Bunny-Phi-2-SigLIP, Bunny-Phi-1.5-SigLIP, Bunny-Phi-2-EVA, and Bunny-Phi-1.5-EVA) [22], CogVLM [69], CogAgent [23], InstructBLIP family (InstructBLIP-Vicuna-7B and InstructBLIP-Vicuna-13B) [13], and LLaVA family (LLaVA-1.6-Mistral-7B, LLaVA-1.6-Vicuna-7B, LLaVA-1.6-Vicuna-13B, and LLaVA-1.6-34B)  [38]. We also evaluate the proprietary models: Open AI’s GPT-4V, GPT-4o, GPT-4, Google Gemini Pro 1.0, and Anthropic Claude 3 Opus.
Evaluation. As each question contains four options, we use accuracy as the main evaluation metric. The same user prompt is appended at the end of each question: First, provide a concise answer in one sentence. Then, elaborate on the reasoning behind your answer in a detailed, step-by-step explanation. For each model, we adopt the default configurations and decoding strategies, e.g., $\mathsf{argmax}$ for deterministic decoding and top-$p$ for non-deterministic decoding. For non-deterministic decoding, the results are averaged over three independent runs for open-source models. For proprietary models, due to their limited availability and increased compute time and cost, we only perform one run. We summarize the terminologies regarding input modalities for LLMs and VLMs in Table 1:
Model Input Modality Term Description
LLM Text-only TQA (LLM) Text-only input that includes all necessary information to answer questions without visual context.
VLM Text-only TQA (VLM) Text-only input as in TQA (LLM) but applied to VLMs (e.g., the LLaVA family).
VLM Vision-only VQA Input only includes an image without corresponding textual description.
VLM Vision-text VTQA Input includes both an image and its textual description.
Table 1: Terms regarding input modalities for LLMs and VLMs.
We describe the Text-only, Vision-only, and Vision-text input modalities based on how we feed the image information to the models. Vision-only input means the image is fed directly to the models without textual description, while all questions are presented in text.
@@ 3 Main Results and Analysis
Figure 5: Performance overview on spatial reasoning tasks. We report the accuracy averaged over all questions. We consider the VQA (Vision-only) format for vision-language models. The dashed red line denotes the expected accuracy for random guessing. For Spatial-Map and Maze-Nav tasks, only a few models outperform random guessing by a notable margin.
Spatial reasoning remains surprisingly challenging. The evaluation results on open-source models on Spatial-Map, Maze-Nav, and Spatial-Grid are shown in Figure 5. For each task, the reported accuracy is averaged over all questions. For vision-language models, we choose the Vision-only input format, commonly used in Visual Question Answering (VQA). We use a dashed red line in each figure to indicate the expected accuracy if answering by random guessing. Our findings reveal several notable insights: (1) Vision-only inputs: Despite the simplicity of these tasks for humans, most competitive multimodal models perform at levels similar to or barely above random guessing. (2) Text-only inputs: While the textual input includes essential spatial information, it generally does not significantly enhance the spatial reasoning capabilities of competitive models. An exception occurs in the Spatial-Grid task, where Llama-3 achieves an accuracy of 71.9%, followed by Mistral-7B-Instruct at 62.1%, both notably surpassing random guessing. Despite these successes, the performance of these models still lags significantly behind human levels. These results underscore the need for further development of techniques tailored to spatial understanding and reasoning.
Figure 6: TQA (LLM) vs. VQA on spatial reasoning tasks. Each vertex on the spider plot represents the Avg Acc of a (VLM, LLM) pair with the same language backbone, i.e., LLM v.s. VLM further finetuned on that. VLMs are depicted in red, and LLMs in blue. We can see that VLMs rarely enhance the performance compared to their LLM counterparts.
The impact of input modality.
To investigate the impact of modality, we compare the performance of a large language model (LLM) and a vision-language model (VLM) with the same language backbone. We consider the VQA (Vision-only) format for vision-language models. The results are shown in Figure 6. Each vertex on the spider plot represents the average accuracy of a (VLM, LLM) pair. We observe that on Spatial-Map and Spatial-Grid, the majority of VLMs yield worse performance compared to their LLM counterpart, despite having an additional visual encoder. For example, on Spatial-Grid, Mixtral-7B achieves an average accuracy of 62.1%, while LLaVA-v1.6-Mistral-7B only yields an accuracy of 47.1% (15% $\downarrow$). Detailed results can be seen in Appendix F.
@@ 4.1 Seeing Without Understanding: The Blindness of Multimodal Language Models
To better understand how VLMs process visual information, we conduct a series of controlled experiments in the VTQA (Vision-text input) setting. For each sample, we replace the original image input (that matches the textual description) with either: (1) No Image: only keep the textual input without the image input, (2) Noise Image: a Gaussian noise image irrelevant to the task, and (3) Random Image: a random image from the dataset that does not match the textual description, as shown in Figure 7.
Figure 8: VTQA vs. TQA (VLM) on spatial reasoning tasks. VLMs exhibit improved performance in spatial reasoning tasks when visual input is absent.
@@
Noise image can improve the performance. We replace the Original Image with a Noise Image while retaining the original textual description. The results are shown in Figure 9. Consistent with the findings in Original Image vs. No Image, using a noise image also improves the performance across various VLM architectures. For example, the accuracy of LLaVA-1.6-Vicuna-13B increases by 6.5% on the Maze-Nav task when the noise image is used as opposed to the original image. In contrast to the No Image setting, noise images provide limited visual cues. Nonetheless, the model tends to prioritize the textual information, especially when visual cues are not pertinent to the task.
@@ 4.2 Leveraging Redundancy in Multimodal Inputs
Multimodal language models offer considerable versatility in handling multimodal inputs. While the visual input alone often provides sufficient details for humans to address spatial reasoning tasks with ease, we propose that VLMs significantly benefit from the inclusion of textual descriptions alongside visual data, even if this introduces substantial redundancy. We verify this hypothesis by comparing VQA (Vision-only input) and VTQA (Vision-text input) across diverse VLM architectures. The results are shown in Figure 11, where each vertex on the spider plot represents the average accuracy of a (Vision-only, Vision-text) pair based on the same VLM. For Spatial-Map and Spatial-Grid, we can clearly see that having the additional textual input (VTQA) enhances the performance compared to only using images (VQA) across different VLM architectures. This suggests that textual inputs improve the accuracy of spatial reasoning in VLMs. We further compare TQA and VTQA in Appendix D. Detailed results are included in Appendix F.
@@
Comparison Results Summary of Findings
TQA (LLM) vs. VQA Figure 6 VQA rarely enhances the performance compared to TQA (LLM).
VTQA vs. TQA (VLM) Figure 8 VLMs exhibit improved performance in spatial reasoning tasks when the image input is absent.
VQA vs. VTQA Figure 11 Given the same image input, additional textual description enhances VLM’s performance.
TQA (VLM) vs. TQA (LLM) Figure 12 Multimodal fine-tuning enhances LLM’s spatial reasoning ability.
TQA (LLM) vs. VTQA Figure 16 No definitive winner.
Table 2: Summary of main findings.
As many recent benchmarks have shown that proprietary models generally outperform open-source models, it is important to understand if our observed trends hold with proprietary models. The performance of several top proprietary models (GPT-4, GPT-4V, GPT-4o, Gemini Pro 1.0, and Claude 3 Opus) are shown in Figure 13. We have the following salient observations: (1) A significant performance gap exists between SoTA open-source models and proprietary models, as expected. Furthermore, with both Text-only and Vision-text formats, GPT-4V and GPT-4o significantly outperform random guessing across all tasks. For instance, in the Vision-text format, GPT-4o achieves an accuracy of 0.989 on Spatial-Grid (Table 7). (2) Yet, the trends we observed with open source models hold, VQA consistently under-performs compared to TQA and VTQA, for example, GPT-4V’s performance improves by 25.6% on Spatial-Grid when switching from Vision-only to Vision-text input; and again no clear winner between TQA and VTQA (see Appendix D for further details), showing that proprietary models, even the new GPT-4o model, still do not appear to fully leverage visual inputs. We summarize the key findings of this work in Table 2.
5 Related Work
Large language models. Large language models (LLMs) have achieved outstanding performance across a diverse array of fields, including finance [34], bioinformatics [63], law [60], education [29], coding [24], and creative tasks [2, 36]. LLM architectures have gone through significant changes in recent years, with notable developments such as BERT [14], OPT [76], PaLM [12], Gemma family [62], Mistral family [27], GPT family [2, 10], Claude family [5], and LLaMA family [3, 64, 66]. These models have demonstrated emergent abilities and revolutionized numerous domains, supporting capabilities such as in-context learning [46, 51, 59], compositional reasoning [16, 20, 70], and task-specific adaptation [57, 58, 71]. However, visual understanding and reasoning—an intrinsic part of human cognitive ability—remains largely under-explored for LLMs.
Vision-language models and multi-modal language models. The success of LLMs has propelled the adoption of the Transformer architecture [68] within the computer vision community, such as ViT [15], Beit [9], CLIP [54], MAE [21], Swin [42, 43], and DiT [53]. Building on the capabilities of powerful LLMs, multi-modal language models (MLLMs) such as Flamingo [4], LLaMA-Adapter [19, 74], LLava [37, 39], stable-diffusion [55], BLIP [32, 33], MiniGPT-4 [77], Qwen [7, 8], Gemini [61], MM1 [45] have significantly expanded the range of problems that can be addressed with improved reliability [49]. These models adeptly handle inputs from diverse modalities and have demonstrated remarkable performance on diverse tasks such as mathematical reasoning [44], image-text retrieval [47, 73], and visual reasoning [6, 18, 25, 28, 31, 40, 50].